Three-circle, red-on-white is one distinctive pattern painted on ceramic vessels of the Anasazi period found at an archaeological site. At one excavation, a sample of 179 potsherds indicated that 78 were of the three-circle, red-on-white pattern.

(a) Find a point estimate p̂ for the proportion of all ceramic potsherds at this site that are of the three-circle, red-on-white pattern. (Round your answer to four decimal places.)

(b) Compute a 95% confidence interval for the population proportion p of all ceramic potsherds with this distinctive pattern found at the site. (Round your answers to three decimal places.)

lower limit

upper limit
Answer #1

a)

point estimate = X/n = 78/179 =

0.435754

b)

95% confidence interval

(point estimate +- margin of error)

margin of error

= z *sqrt(pq/n) {z = 1.96 for 95%}

0.072641

hence

lower 0.363113
upper 0.508396